OpenStudy (anonymous):

how do you set up this? a car radiator contains 10 liters of 30% solution. how many liter will have to be replaced with pure anti-freeze if the resulting solution is to be 50%?

hero (hero):

x = amount of pure antifreeze y = total radiator fluid solution Total liters of Radiator Fluid Solution: 10 + x = y Total Percent Liters of Radiator Fluid Solution: 10(.30) + x = .50y Solve this system: 10 + x = y 10(.30) + x = .50y

OpenStudy (anonymous):

THANKS!!!

hero (hero):

What did you get for the values of x and y?